# Hierarchical Time Series Forecasting with Robust Reconciliation

### v1

@@ Abstract
This paper focuses on forecasting hierarchical time-series data, where each higher-level observation equals the sum of its corresponding lower-level time series. In such contexts, the forecast values should be coherent, meaning that the forecast value of each parent series exactly matches the sum of the forecast values of its child series. Existing hierarchical forecasting methods typically generate base forecasts independently for each series and then apply a reconciliation procedure to adjust them so that the resulting forecast values are coherent across the hierarchy. These methods generally derive an optimal reconciliation, using a covariance matrix of the forecast error. In practice, however, the true covariance matrix is unknown and has to be estimated from finite samples in advance. This gap between the true and estimated covariance matrix may degrade forecast performance. To address this issue, we propose a robust optimization framework for hierarchical reconciliation that accounts for uncertainty in the estimated covariance matrix. We first introduce an uncertainty set for the estimated covariance matrix and formulate a reconciliation problem that minimizes the worst-case expected squared error over this uncertainty set. We show that our problem can be cast as a semidefinite optimization problem. Numerical experiments demonstrate that the proposed robust reconciliation method achieved better forecast performance than existing hierarchical forecasting methods, which indicates the effectiveness of integrating uncertainty into the reconciliation process.
@@ 1 Introduction
Time series forecasting is indispensable across diverse fields, including sales planning and inventory management (Aviv, 2003; Ramos et al., 2015), energy supply planning (Suganthi and Samuel, 2012; Hernandez et al., 2014), and economic analysis and stock investment decision-making (Krollner et al., 2010). For instance, in the retail sector, accurate sales predictions based on historical data are crucial for optimizing inventory levels and preventing both overstocking and shortages. Similarly, for electric power companies, forecasting electricity consumption enables efficient facility operation and effective supply-demand balance management. Moreover, at both individual and national levels, leveraging forecasts of economic indicators and stock prices can contribute significantly to wealth generation. Conversely, low forecast accuracy can lead to substantial losses and missed opportunities for individuals, businesses, and society as a whole. Consequently, extensive research has been dedicated to developing various time series forecasting methods, and there remains a strong demand for more precise techniques (Mahalakshmi et al., 2016; Liu et al., 2021; Wen et al., 2022).
@@
Many real-world datasets inherently possess hierarchical structures. Examples include sales data organized by region or demographic statistics categorized by gender or age group, which are commonly recorded across multiple levels of aggregation. In practice, the appropriate level of the hierarchy for forecasting depends on the specific application, and this choice can significantly influence the prediction outcomes. Generally, as one descends to lower hierarchical levels, the data becomes more granular but also more susceptible to individual variations and noise, leading to increased uncertainty. This often means that aggregated data at higher levels tends to be more stable and achieve greater forecast accuracy (Grunfeld and Griliches, 1960). However, higher-level aggregated data can obscure fine-grained patterns and individual variation factors. Therefore, it has also been suggested that utilizing detailed data from lower levels, if appropriately modeled, can potentially yield superior forecast accuracy (Orcutt et al., 1968; Edwards and Orcutt, 1969).
@@
Given this context, time series forecasting methods that explicitly account for hierarchical structures have garnered increasing attention (Athanasopoulos et al., 2009; Hyndman et al., 2011; Wickramasuriya et al., 2019; Panagiotelis et al., 2021; Hyndman and Athanasopoulos, 2021). These approaches aim to adjust forecasts across both lower and higher levels to ensure coherence when aggregating forecast values within the hierarchy. Such methods are expected to enhance forecast accuracy compared to conventional time series forecasting based on a single level.
@@
Despite these advancements, existing hierarchical time series forecasting methods face certain challenges. Traditional approaches typically aim to minimize the expected forecast error at a given target time point, which necessitates a covariance matrix of the forecast errors. This matrix is commonly estimated from the residuals between observed and forecast values. However, if the underlying data trends shift or the forecasting model is not sufficiently accurate, discrepancies can emerge between the estimated and true covariance matrices. Thus, the estimated covariance matrix itself carries inherent uncertainty, which must be addressed. Prior work by Møller et al. (2024) focused on this issue, decomposing covariance matrix estimation into parameter estimation errors and stochastic irreducible errors to quantify uncertainty and improve forecast accuracy. Nevertheless, even with their method, the true covariance matrix cannot be perfectly determined, leaving room for further improvements in forecast accuracy.
@@
To tackle the uncertainty of estimators, robust optimization has emerged as a powerful technique. This methodology is designed to yield solutions that remain effective even when the underlying data fluctuates within a defined uncertain range. Specifically, it involves establishing a range for uncertain parameters or data and then seeking an optimal solution that performs best under the worst-case scenario within that range. Since its inception by Ben-Tal and Nemirovski (1998), robust optimization has been extensively researched in both theoretical and applied domains (Bertsimas et al., 2011). Notably, models that incorporate covariance matrix uncertainty have been developed and applied to various problems, such as portfolio optimization (Lobo and Boyd, 2000; Halldórsson and Tütüncü, 2003; Tütüncü and Koenig, 2004).
@@
In this paper, we propose a novel method that frames hierarchical time series forecasting as a robust optimization problem, specifically minimizing forecast error under uncertainty. We introduce an uncertainty set for the covariance matrix of forecast errors and solve an optimization problem that minimizes the forecast error under the worst-case scenario within this set. We demonstrate through duality that this robust optimization problem can be formulated as a semidefinite optimization problem, which is theoretically solvable efficiently. Furthermore, we present numerical experiments using five real-world datasets, showcasing that our proposed method achieves more accurate forecasting results compared to existing hierarchical time series forecasting techniques.
2 Hierarchical time series forecasting
@@ 2.1 Notation
This section defines the notation used throughout the paper, which is consistent with prior studies on hierarchical time series forecasting (Athanasopoulos et al., 2009; Hyndman et al., 2011; Wickramasuriya et al., 2019; Panagiotelis et al., 2021; Hyndman and Athanasopoulos, 2021).
A hierarchical structure is defined by a series of nested levels. Level 0 is the fully aggregated series. Level 1 consists of the series obtained by disaggregating the Level 0 series, and Level 2 contains the series that further disaggregate each Level 1 series. This process continues until the bottom-level, denoted as Level $K$, where its series can no longer be disaggregated.
@@
When Equation˜1 holds for the values of all series and the bottom-level series at each time $t$, it is said the hierarchy is satisfied.
@@ 2.2 Reconciliation methods
Hierarchical time series forecasting is a process that adjusts, or “reconciles,” forecast values to ensure they satisfy the hierarchical aggregation constraints. We assume that observations for each series are available for an observation period $t=1,\ldots,T$, and the objective is to forecast values for the forecast period $\tau=T+1,\ldots,T+T^{\prime}$.
Forecasting, ignoring any aggregation constraints, is called a base forecast. Let $\hat{\bm{y}}^{\left(\tau\right)}\in\mathbb{R}^{n}$ denote the vector of base forecasts for all series at time $\tau$. Forecasting that satisfies the hierarchical aggregation constraints is called a coherent forecast. Let $\tilde{\bm{y}}^{\left(\tau\right)}\in\mathbb{R}^{n}$ denote the vector of coherent forecasts at time $\tau$. To transform base forecasts into coherent forecasts, hierarchical forecasting methods estimate a reconciliation matrix $\bm{P}\in\mathbb{R}^{m\times n}$ that maps the base forecasts to the bottom-level. All hierarchical forecasting approaches can be expressed in the general form:
$\tilde{\bm{y}}^{\left(\tau\right)}=\bm{S}\bm{P}\hat{\bm{y}}^{\left(\tau\right)},$ (2)
@@
Because the bottom-up and top-down methods utilize base forecasts from only a single level of aggregation, they rely on limited information. To overcome this, subsequent research has proposed methods that use base forecasts from all series to estimate a reconciliation matrix, thus producing more comprehensive coherent forecasts.
@@
Hyndman et al. (2011) proposed the generalized least squares (GLS) reconciliation, a regression-based approach. In this method, the reconciliation matrix is chosen to minimize errors between the base forecasts and the coherent forecasts. Specifically, consider the regression model for base forecasts
$\hat{\bm{y}}^{\left(\tau\right)}=\bm{S}\bm{\beta}^{\left(\tau\right)}+\bm{\varepsilon}^{\left(\tau\right)}.$
@@
Let $\bm{\beta}^{\left(\tau\right)}=\mathbb{E}\left[\bm{b}^{\left(\tau\right)}\mid\bm{y}^{\left(1\right)},\ldots,\bm{y}^{\left(T\right)}\right]\in\mathbb{R}^{m}$ be the expectation of base forecasts at the bottom-level, and let the error term $\bm{\varepsilon}^{\left(\tau\right)}$ have zero mean with covariance matrix $\bm{\Sigma}^{\left(\tau\right)}=\operatorname{Var}\left(\bm{\varepsilon}^{\left(\tau\right)}\mid\bm{y}^{\left(1\right)},\ldots,\bm{y}^{\left(T\right)}\right)\in\mathbb{R}^{n\times n}$. If $\bm{\Sigma}^{\left(\tau\right)}$ were known, the minimum variance unbiased estimator of $\bm{\beta}^{\left(\tau\right)}$ would be the GLS estimator
$\hat{\bm{\beta}}^{\left(\tau\right)}=\left(\bm{S}^{\top}\bm{\Sigma}^{\left(\tau\right)\dagger}\bm{S}\right)^{-1}\bm{S}^{\top}\bm{\Sigma}^{\left(\tau\right)\dagger}\hat{\bm{y}}^{\left(\tau\right)},$
@@
where $\bm{\Sigma}^{\left(\tau\right)\dagger}$ is the generalized inverse of $\bm{\Sigma}^{\left(\tau\right)}$. Comparing this expression with Equation˜2 gives
$\bm{P}=\left(\bm{S}^{\top}\bm{\Sigma}^{\left(\tau\right)\dagger}\bm{S}\right)^{-1}\bm{S}^{\top}\bm{\Sigma}^{\left(\tau\right)\dagger}.$
@@
In practice, however, the covariance matrix $\bm{\Sigma}^{\left(\tau\right)}$ is unknown and cannot be estimated. It represents the covariance of the reconciliation errors at time $\tau$, but the errors are not observable until coherent forecasts have already been produced. Hyndman et al. (2011) shows that, under the strong assumption that the errors themselves satisfy the aggregation constraints, $\bm{\Sigma}^{\left(\tau\right)}$ can be replaced with the $n$-dimensional identity matrix $\bm{I}_{n}$. This replacement is equivalent to computing the ordinary least squares (OLS) estimator instead of the GLS estimator.
@@
Wickramasuriya et al. (2019) proposed the minimum-trace (MinT) reconciliation, which does not rely on the assumptions required by Hyndman et al. (2011). MinT chooses the reconciliation matrix to minimize the variance of errors between the observation values and the coherent forecasts. Let
$\bm{V}^{\left(\tau\right)}=\operatorname{Var}\left(\bm{y}^{\left(\tau\right)}-\tilde{\bm{y}}^{\left(\tau\right)}\mid\bm{y}^{\left(1\right)},\ldots,\bm{y}^{\left(T\right)}\right)\in\mathbb{R}^{n\times n}$
@@
denote the covariance matrix of those errors and define the covariance matrix of the errors between observation values and base forecasts as
$\bm{W}^{\left(\tau\right)}=\operatorname{Var}\left(\bm{y}^{\left(\tau\right)}-\hat{\bm{y}}^{\left(\tau\right)}\mid\bm{y}^{\left(1\right)},\ldots,\bm{y}^{\left(T\right)}\right)\in\mathbb{R}^{n\times n}.$
With these matrices, we can derive the relationship as follows:
$\bm{V}^{\left(\tau\right)}=\bm{S}\bm{P}\bm{W}^{\left(\tau\right)}\bm{P}^{\top}\bm{S}^{\top}.$
@@
Then the optimal reconciliation matrix, which minimizes the total error variance $\operatorname{tr}\left(\bm{V}^{\left(\tau\right)}\right)$ is given by
$\bm{P}=\left(\bm{S}^{\top}\left(\bm{W}^{\left(\tau\right)}\right)^{-1}\bm{S}\right)^{-1}\bm{S}^{\top}\left(\bm{W}^{\left(\tau\right)}\right)^{-1}.$
Although the covariance matrix $\bm{W}^{\left(\tau\right)}$ is unknown, it can be estimated using the unbiased sample variance calculated from the observations and the base forecasts at times $t=1,\ldots,T$. Because the same estimate is applied to every forecast period $\tau=T+1,\ldots,T+T^{\prime}$, we denote it by $\bm{W}$ and set $\bm{W}^{\left(\tau\right)}=\bm{W}$ for all $\tau$. The covariance matrix is always positive semidefinite but not always positive definite, and when it is not positive definite, it cannot be inverted. A common remedy is to apply a shrinkage approach (Schäfer and Strimmer, 2005) to obtain a well-conditioned matrix instead of the original covariance matrix.
Wickramasuriya et al. (2019) proposed MinT as an approach that minimizes the variance of the errors between the observation values and the coherent forecasts, while Panagiotelis et al. (2021) showed that MinT is equivalent to minimizing the expected value of the errors. Rewriting $\operatorname{tr}\left(\bm{V}^{\left(\tau\right)}\right)$ gives
$\operatorname{tr}\left(\bm{V}^{\left(\tau\right)}\right)$ $=\operatorname{tr}\left(\mathbb{E}\left[\left(\bm{y}^{\left(\tau\right)}-\tilde{\bm{y}}^{\left(\tau\right)}\right)\bm{W}\left(\bm{y}^{\left(\tau\right)}-\tilde{\bm{y}}^{\left(\tau\right)}\right)^{\top}\right]\right)$
$=\mathbb{E}\left[\left(\bm{y}^{\left(\tau\right)}-\tilde{\bm{y}}^{\left(\tau\right)}\right)^{\top}\bm{W}\left(\bm{y}^{\left(\tau\right)}-\tilde{\bm{y}}^{\left(\tau\right)}\right)\right].$
Consequently, computing the reconciliation matrix by minimizing $\operatorname{tr}\left(\bm{V}^{\left(\tau\right)}\right)$ is equivalent to solving the optimization problem
$\min_{\bm{P}}\quad$ $\mathbb{E}\left[\left(\bm{y}^{\left(\tau\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(\tau\right)}\right)^{\top}\bm{W}\left(\bm{y}^{\left(\tau\right)}-\bm{S}\bm{P}\hat{\bm{y}}^{\left(\tau\right)}\right)\right].$ (3)

### v2

@@ Abstract
This paper focuses on forecasting hierarchical time-series data, where each higher-level observation equals the sum of its corresponding lower-level time series. In such contexts, the forecast values should be coherent, meaning that the forecast value of each parent series exactly matches the sum of the forecast values of its child series. Existing hierarchical forecasting methods typically generate base forecasts independently for each series and then apply a reconciliation procedure to adjust them so that the resulting forecast values are coherent across the hierarchy. These methods generally yield an optimal reconciliation, using a covariance matrix of the forecast errors. In practice, however, the true covariance matrix is unknown and has to be estimated from finite samples in advance. This gap between the true and estimated covariance matrix may degrade forecast performance. To address this issue, we propose a robust optimization framework for hierarchical reconciliation that accounts for uncertainty in the estimated covariance matrix. We first introduce an uncertainty set for the estimated covariance matrix and formulate a reconciliation problem that minimizes the worst-case average of weighted squared residuals over this uncertainty set. We show that our problem can be cast as a semidefinite optimization problem. Numerical experiments demonstrate that the proposed robust reconciliation method achieved better forecast performance than existing hierarchical forecasting methods, which indicates the effectiveness of integrating uncertainty into the reconciliation process.
@@ 1 Introduction
Time-series forecasting is indispensable across diverse fields, including sales planning and inventory management (Aviv, 2003; Ramos et al., 2015), energy supply planning (Suganthi and Samuel, 2012; Hernandez et al., 2014), and economic analysis and stock investment decision-making (Krollner et al., 2010). For instance, in the retail sector, accurate sales predictions based on historical data are crucial for optimizing inventory levels and preventing both overstocking and shortages. Similarly, for electric power companies, forecasting electricity consumption enables efficient facility operations and effective supply-demand balance management. Moreover, at both the individual and national levels, leveraging forecasts of economic indicators and stock prices can significantly contribute to wealth creation. Conversely, low forecast accuracy can lead to substantial losses and missed opportunities for individuals, businesses, and society as a whole. Consequently, extensive research has focused on developing various time-series forecasting methods, and there remains a strong demand for more precise techniques (Mahalakshmi et al., 2016; Liu et al., 2021; Wen et al., 2022).
@@
Many real-world datasets inherently possess hierarchical structures. Examples include sales data organized by region or demographic statistics categorized by gender or age group, which are commonly recorded across multiple levels of aggregation. In practice, the appropriate hierarchical level for forecasting depends on the specific application, and this choice can significantly influence prediction outcomes. Generally, as one descends to lower hierarchical levels, the data becomes more granular but also more susceptible to individual variations and noise, leading to increased uncertainty. This often means that aggregated data at higher levels tends to be more stable and achieve greater forecast accuracy (Grunfeld and Griliches, 1960). However, higher-level aggregated data can obscure fine-grained patterns and individual variation factors. Therefore, it has also been suggested that utilizing detailed data from lower levels, if appropriately modeled, can potentially yield superior forecast accuracy (Orcutt et al., 1968; Edwards and Orcutt, 1969).
@@
Given this context, time-series forecasting methods that explicitly account for hierarchical structures have garnered increasing attention (Athanasopoulos et al., 2009; Hyndman et al., 2011; Wickramasuriya et al., 2019; Shiratori et al., 2020; Hyndman and Athanasopoulos, 2021). These approaches aim to adjust forecasts across both lower and higher levels to ensure coherence when aggregating forecast values within the hierarchy. Such methods are expected to enhance forecast accuracy compared to conventional time-series forecasting based on a single level.
@@
Despite these advancements, existing hierarchical time-series forecasting methods face certain challenges. Traditional approaches typically aim to minimize the expected forecast error at a given target time point, which necessitates a covariance matrix of the forecast errors. This matrix is commonly estimated from the residuals between observed and forecast values. However, if underlying data trends shift or the forecasting model is inaccurate, discrepancies can arise between the estimated and true covariance matrices. Thus, the estimated covariance matrix itself carries inherent uncertainty, which must be addressed. Prior work by Møller et al. (2024) focused on this issue, decomposing covariance matrix estimation into parameter estimation errors and stochastic irreducible errors to quantify uncertainty and improve forecast accuracy. Nevertheless, even with their method, the true covariance matrix cannot be perfectly determined, leaving room for further improvements in forecast accuracy.
@@
To address the uncertainty inherent in estimators, robust optimization has emerged as a powerful technique. This methodology is designed to yield solutions that remain effective even when the underlying data fluctuates within a defined uncertain range. Specifically, it involves establishing a range for uncertain parameters or data and then seeking an optimal solution that performs best under the worst-case scenario within that range. Since its inception by Ben-Tal and Nemirovski (1998), robust optimization has been extensively researched in both theoretical and applied domains (Bertsimas et al., 2011). Notably, models that incorporate covariance matrix uncertainty have been developed and applied to various problems, such as portfolio optimization (Lobo and Boyd, 2000; Halldórsson and Tütüncü, 2003; Tütüncü and Koenig, 2004).
@@
In this paper, we propose a novel method that frames hierarchical time-series forecasting as a robust optimization problem. Our approach aims to minimize the average of forecast residuals over an observation period under the worst-case scenario within an uncertainty set for the covariance matrix of forecast errors. We demonstrate through duality that this robust optimization problem can be formulated as a semidefinite optimization problem, which is theoretically solvable efficiently. Furthermore, we present numerical experiments on five real-world datasets, showing that our proposed method achieves more accurate forecasts than existing hierarchical time-series forecasting techniques.
2 Hierarchical time-series forecasting
@@ 2.1 Notation
This section defines the notation used throughout the paper, which is consistent with prior studies on hierarchical time-series forecasting (Athanasopoulos et al., 2009; Hyndman et al., 2011; Wickramasuriya et al., 2019; Panagiotelis et al., 2021; Hyndman and Athanasopoulos, 2021).
A hierarchical structure is defined by a series of nested levels. Level 0 is the fully aggregated series. Level 1 consists of the series obtained by disaggregating the Level 0 series, and Level 2 contains the series that further disaggregate each Level 1 series. This process continues until the bottom level, denoted as Level $K$, where its series can no longer be disaggregated.
@@
When Equation˜1 holds for the values of all series and the bottom-level series at each time $t$, it is said that the hierarchy is satisfied.
@@ 2.2 Reconciliation methods
Hierarchical time-series forecasting is a process that adjusts, or “reconciles,” forecast values to ensure they satisfy the hierarchical aggregation constraints. We assume that observations for each series are available for an observation period $t=1,\ldots,T$, and the objective is to forecast values for the forecast period $\tau=T+1,\ldots,T+T^{\prime}$.
Forecasting, ignoring any aggregation constraints, is called a base forecast. Let $\hat{\bm{y}}^{(\tau)}\in\mathbb{R}^{n}$ denote the vector of base forecasts for all series at time $\tau$. A forecast that satisfies the hierarchical aggregation constraints is called a coherent forecast. Let $\tilde{\bm{y}}^{(\tau)}\in\mathbb{R}^{n}$ denote the vector of coherent forecasts at time $\tau$. To transform base forecasts into coherent forecasts, hierarchical forecasting methods estimate a reconciliation matrix $\bm{P}\in\mathbb{R}^{m\times n}$ that maps the base forecasts to the bottom level. All hierarchical forecasting approaches can be expressed in the general form:
$\tilde{\bm{y}}^{(\tau)}=\bm{S}\bm{P}\hat{\bm{y}}^{(\tau)},$ (2)
@@
Because the bottom-up and top-down methods utilize base forecasts from only a single level of aggregation, they rely on limited information. To overcome this, subsequent research has proposed methods that use base forecasts from all series to estimate a reconciliation matrix, thus producing more comprehensive, coherent forecasts.
@@
Hyndman et al. (2011) proposed the generalized least squares (GLS) reconciliation, a regression-based approach. In this method, the reconciliation matrix is chosen to minimize errors between the base and coherent forecasts. Specifically, consider the regression model for base forecasts
$\hat{\bm{y}}^{(\tau)}=\bm{S}\bm{\beta}^{(\tau)}+\bm{\varepsilon}^{(\tau)}.$
@@
Let $\bm{\beta}^{(\tau)}=\mathbb{E}\left[\bm{b}^{(\tau)}\mid\bm{y}^{\left(1\right)},\ldots,\bm{y}^{\left(T\right)}\right]\in\mathbb{R}^{m}$ be the expectation of base forecasts at the bottom level, and let the error term $\bm{\varepsilon}^{(\tau)}$ have zero mean with covariance matrix $\bm{\Sigma}^{(\tau)}=\operatorname{Var}\left(\bm{\varepsilon}^{(\tau)}\mid\bm{y}^{\left(1\right)},\ldots,\bm{y}^{\left(T\right)}\right)\in\mathbb{R}^{n\times n}$. If $\bm{\Sigma}^{(\tau)}$ were known, the minimum variance unbiased estimator of $\bm{\beta}^{(\tau)}$ would be the GLS estimator
$\hat{\bm{\beta}}^{(\tau)}=\left(\bm{S}^{\top}\bm{\Sigma}^{(\tau)\dagger}\bm{S}\right)^{-1}\bm{S}^{\top}\bm{\Sigma}^{(\tau)\dagger}\hat{\bm{y}}^{(\tau)},$
@@
where $\bm{\Sigma}^{(\tau)\dagger}$ is the generalized inverse of $\bm{\Sigma}^{(\tau)}$. Comparing this expression with Equation˜2 gives
$\bm{P}=\left(\bm{S}^{\top}\bm{\Sigma}^{(\tau)\dagger}\bm{S}\right)^{-1}\bm{S}^{\top}\bm{\Sigma}^{(\tau)\dagger}.$
@@
In practice, however, the covariance matrix $\bm{\Sigma}^{(\tau)}$ is unknown and cannot be estimated. It represents the covariance of the reconciliation errors at time $\tau$, but the errors are not observable until coherent forecasts have already been produced. Hyndman et al. (2011) showed that, under the assumption that the errors themselves satisfy the aggregation constraints, $\bm{\Sigma}^{(\tau)}$ can be replaced with the $n$-dimensional identity matrix $\bm{I}_{n}$. This replacement is equivalent to computing the ordinary least squares (OLS) estimator instead of the GLS estimator.
@@
Wickramasuriya et al. (2019) proposed the minimum-trace (MinT) reconciliation, which determines the reconciliation matrix $\bm{P}$ by minimizing the total variance of errors between the observed values and the coherent forecasts. Let
$\bm{V}^{(\tau)}=\operatorname{Var}\left(\bm{y}^{(\tau)}-\tilde{\bm{y}}^{(\tau)}\mid\bm{y}^{\left(1\right)},\ldots,\bm{y}^{\left(T\right)}\right)\in\mathbb{R}^{n\times n}$
@@
denote the covariance matrix of those errors, and define the covariance matrix of the errors between observed values and base forecasts as
$\bm{W}^{(\tau)}=\operatorname{Var}\left(\bm{y}^{(\tau)}-\hat{\bm{y}}^{(\tau)}\mid\bm{y}^{\left(1\right)},\ldots,\bm{y}^{\left(T\right)}\right)\in\mathbb{R}^{n\times n}.$
@@
Under the linear transformation in Equation˜2, the error covariance matrix of the coherent forecasts is expressed as $\bm{V}^{(\tau)}=\bm{S}\bm{P}\bm{W}^{(\tau)}\bm{P}^{\top}\bm{S}^{\top}$. Consequently, the optimal reconciliation matrix that minimizes the total error variance $\operatorname{tr}\left(\bm{V}^{(\tau)}\right)$ is given by
$\bm{P}=\left(\bm{S}^{\top}\left(\bm{W}^{(\tau)}\right)^{-1}\bm{S}\right)^{-1}\bm{S}^{\top}\left(\bm{W}^{(\tau)}\right)^{-1}.$ (3)
A notable property of this reconciliation is that the resulting coherent forecasts $\tilde{\bm{y}}^{(\tau)}$ are guaranteed to be at least as accurate as the base forecasts $\hat{\bm{y}}^{(\tau)}$ in terms of the weighted squared error using $\left(\bm{W}^{(\tau)}\right)^{-1}$:
$\left(\bm{y}^{(\tau)}-\tilde{\bm{y}}^{(\tau)}\right)^{\top}\left(\bm{W}^{(\tau)}\right)^{-1}\left(\bm{y}^{(\tau)}-\tilde{\bm{y}}^{(\tau)}\right)\leq\left(\bm{y}^{(\tau)}-\hat{\bm{y}}^{(\tau)}\right)^{\top}\left(\bm{W}^{(\tau)}\right)^{-1}\left(\bm{y}^{(\tau)}-\hat{\bm{y}}^{(\tau)}\right).$
As noted by Panagiotelis et al. (2021), when the base forecast $\hat{\bm{y}}^{(\tau)}$ is unbiased, the trace minimization in MinT is equivalent to minimizing the expected squared Euclidean norm of the errors between the observed values and the coherent forecasts:
$\min_{\bm{P}}\quad\mathbb{E}[\lVert\bm{y}^{(\tau)}-\bm{S}\bm{P}\hat{\bm{y}}^{(\tau)}\rVert^{2}],$ (4)
where $\lVert\cdot\rVert$ is the Euclidean norm. Furthermore, they provided a key insight that extends this Euclidean norm to a more general weighted norm. Specifically, they showed that if we know the true covariance matrix $\bm{W}^{(\tau)}$, the solution in Equation˜3 is invariant to the choice of the weight matrix $\bm{M}$ in the following generalized optimization problem:
$\min_{\bm{P}}\quad\mathbb{E}\left[(\bm{y}^{(\tau)}-\bm{S}\bm{P}\hat{\bm{y}}^{(\tau)})^{\top}\bm{M}(\bm{y}^{(\tau)}-\bm{S}\bm{P}\hat{\bm{y}}^{(\tau)})\right],$ (5)
where $\bm{M}$ is any positive definite matrix. This property implies that the specific weighting of errors does not alter the optimal reconciliation matrix $\bm{P}$.